# Detect duality obstruction of calibrations in smooth category

### v2

@@ Abstract.
Calibration, duality obstruction, realization problem
@@ 1. Introduction
In [Z12, z] the author developed a very “soft” framework about extending calibration pairs which play effective roles in detecting duality obstruction and realizing problem about singularities (to be explained in below). By combining these methods the framework has been carried forward to attack several long-term standing conjectures regarding area-minimizing currents, e.g. [L1, L2, L3].
@@
This obstruction in $(\ast)$ comes from some delicate balance between geometry and topology. And it has been further pointed out on page 431 of [FH] that our obstruction even exists in local around the submaniold in the category of continuous calibration forms. To distinguish this local obstruction $(\ast)$ from global ones, the author used the following example to explain to others since 2020 and it should be faithfully mentioned that such kind of illustration can trace back at least to [FH]. Let $Y=S^{1}\times S^{2020}$ and choose a closed embedded curve $\gamma_{2}$ to be a representative of twice the generator $\tt e$} of $H_{1}(Y;\mathbb{Z})$. Then as in [Z12] we can conformally change background metric such that $\gamma_{2}$ becomes the unique length minimizer in its homology class. As a result, any length minimizer $\gamma_{1}$ in class $\tt e$} cannot be calibrated by any smooth calibration since otherwise the integral current $2\gamma_{1}$ is also calibrated and thus length minimizing in $\gamma_{2}]$ as well which contradicts with the uniqueness in the above. Here the obstruction comes from global in the sense that one cannot apply the idea in a sufficiently small neighborhood of $\gamma_{2}$ for the duality obstruction phenomenon.
In this paper, we extend our (localizable) duality obstruction for question $(\star)$ to genuine high codimension situations beyond that in $(\ast)$. The idea is to find minimizing cones which support coflat calibrations singular only at the origins but cannot be calibrated by any smooth calibration form, and then apply our extension result for calibration pairs in [Z12] (Theorem 3.1 in below).
@@ 2. Basics of calibrations
The dual complex of smooth forms gives us currents. To be more preciese, by $(\mathscr{E}^{\prime}_{*}(X),d)$ we mean the dual complex of the de Rham complex of $X$. Elements of $\mathscr{E}^{\prime}_{m}(X)$ are $m$-dimensional de Rham currents (with compact support) and $d$ is the adjoint of exterior differentiation.
@@
A beautiful canonical decomposition lemma for an $m$-form with respect to certain $m$-plane in $\mathbb{R}^{n}$ is the following (its proof in [HL1] inspires a remark for $m=1$ and our Lemma 2.12, so we include it in full details).
@@
by $\eta\longmapsto\eta\llcorner\phi,$ namely, $\lambda(\eta)(v)=\phi(\eta\wedge v)$. Set $\eta_{i}=(-1)^{m+i}v_{1}\wedge\cdots\wedge\hat{v}_{i}\wedge\cdots\wedge v_{m}$ for $i=1,\cdots,m$. Consider forms $\lambda_{i}=\lambda(\eta_{i})$. Then we have
@@ 3. New gluing result for the extension of calibration pairs
where $\tt t$ is a nonzero real number ($|t|\geq 1$ since $\|\phi\|^{*}_{g_{E}}\leq 1$). Hence,
@@
As $\phi$ is not zero, there exists $Q$ which realizes $\|\phi\|^{*2}_{g(s)}=\frac{1}{\|Q\|^{2}}$. Then, by dividing $\tt t^{2}$ on both sides of (3.3) and Lemma 2.2, we get
${\|\phi\|^{*2}_{g(s)}}\leq\frac{1-s}{T(Q;0)}+\frac{s}{T(Q;1)}\leq(1-s)\|\phi\|^{*2}_{g_{1}}+s\|\phi\|^{*2}_{g_{2}}.$
@@ 4. Every Lawlor cone supports calibrations singular only at the origin
In this section, we show that every Lawlor cone can support coflat calibrations which are singular only at the origin of the ambient Euclidean space $(\mathbb{R}^{N+1},g_{E})$. Here by Lawlor cone we mean a minimal cone which can be shown area-minimizing according to Lawlor’s curvature criterion (in practice by the c-control or F-control in below). Such a cone naturally supports non-continuous dual calibrations. Based on some analysis in [z], we are able to construct some coflat calibrations resolving the discontinuity and making the singular set exactly be $\{o\}$. Besides, a shortcut for the realization result in [z] will be explained.
@@
In [Law] Lawlor established a similar structure in some suitable angular neighborhood of a minimal cone $C(L^{k})$ instead of in the whole space $\mathbb{R}^{N+1}$. If the boundary of the neighborhood can be mapped to the origin by the projection, then one can map everything outside the neighborhood to the origin. Thus an area-nonincreasing projection is gained and the area-minimality of the cone is proved.
@@
As a result, each Lawlor cone must have some solution $h$ to (4.2) with equality everywhere which touches the $t$-axis at some $t=\tan\theta_{3}>\tan\theta_{0}$ as in the following graph.
Proof of Theorem 1.1. For point $(t,y)$ for $0<y<\sqrt{t^{2}+1}p(t)$, the inequality (4.2) leads to
@@
Note that $h_{0}$ is strictly deceasing, that is to say $h(t)<p(t)$ in the interval $(0,\tan\theta_{0})$. So, one can arrange $a$ and $\delta$ to make $h_{1}$ strictly decreasing as well. Now one gets $h_{1}(t)<p(t)$ and around the graph of $h_{1}$ in the figure the inequality (4.4) tells that the slope $h^{\prime}(t)$ of a solution to the inequality of (4.2) is allowed to vary from a strictly negative number to a strictly positive number as illustrated in the figure. Then one can deform $h_{1}$ to $h_{2}$ which smoothly joins $1-at^{2}$ and the $t$-axis with its graph touching the axis at $\tan\theta_{2}$ as close as one wishes to the point $\tan\theta_{1}$, and moreover satisfies the calibration inequality (4.2). The reason is that in any small interval $\delta,\epsilon+\delta)$ one can extend $1-at^{2}$ smoothly and make the extension fulfill (4.5) near $t=\epsilon+\delta$. This allows us to further extend it by (4.5) and it will hit the $t$-axis at a point $\tan\hat{\theta}$ close to $\tan\theta_{1}$. Similarly as what has been done around $t=\delta$, one can deform it around $\tan\hat{\theta}$ to make it smoothly tangent to the $t$-axis at $\tan\theta_{2}$ such that the resulting $h_{2}$ satisfies (4.2) and $\tan\theta_{2}$ can be as close to $\tan\hat{\theta}$ as one wants.
@@
Based on the observation, we know that $C(L_{1}\times\cdots\times L_{n})$ must be calibrated by $\phi_{o}$ which is a calibration with constant coefficients. Fix $x_{i}\in L_{i}$ and let $\xi_{i}$ be the oriented unit tangent $k_{i}$-vector of $T_{x_{i}}L_{i}$ for $i=2,\cdots,n$. Then we shall focus on the calibration $(k_{1}+1)$-form $\Phi_{o}\triangleq\left(\xi_{2}\wedge\cdots\wedge\xi_{n}\right)\llcorner\phi_{o}$ with constant coefficients. Denote $\mathcal{P}^{*}(\Phi_{o})$ to be the pull-back of $\Phi_{o}$ to the linear subspace $\mathbb{R}^{N_{1}+1}$ via the inclusion $\mathcal{P}:\mathbb{R}^{N_{1}+1}\hookrightarrow\mathbb{R}^{N}$ by $\mathcal{P}(x)=(x,0,\cdots,0)$. Note that for $x\in L_{1}$ we have the oriented unit tangent $k_{1}$-vector $\xi=\xi(x)$ of $T_{x}L_{1}$ and $\xi\wedge x$ is the volume form of $C(L_{1})\subset\mathbb{R}^{N_{1}+1}$ (reverse the orientation if need be, here and in the sequel).
Without loss of generality, assume that $L_{1}$ is connected (or restrict ourselves to one connected component of $L_{1}$). We aim to show that evaluation of $\mathcal{P}^{*}(\Phi_{o})$ with $\xi\wedge x$ is a common nonzero constant for all $x\in L_{1}$. First note that evaluation of $\mathcal{P}^{*}(\Phi_{0})$ with $\xi\wedge x$ has the same evaluation value of $\Phi_{o}$ with $\xi\wedge x$. Moreover, $\Phi_{o}$ (or $\Phi_{o}$) calibrates all the $(k_{1}+1)$-vectors $\{\xi\wedge p\}$ where $p=p(x)=(\lambda_{1}x,\lambda_{2}x_{2},\cdots,\lambda_{n}x_{n})$ with $x$ running in $L_{1}$. So, for each $x\in L_{1}$, we can extend an orthonormal basis $\{v_{1},\cdots,v_{k_{1}}\}$ of span $\xi$ to an orthonormal basis
@@
(5.4) $\left(\mathcal{P}^{*}(\Phi_{0})\right)(\xi\wedge x)$ $=$ $\Phi_{o}(\xi\wedge x)$
@@
By the assumption, we know that $k_{1}+1=N_{1}$. So $\mathcal{P}(\Phi_{0})$ is a simple form in $\mathbb{R}^{N_{1}+1}$. In particular, by (5.4) and the connectedness of $L_{1}$ (or its connected component), the unit normal dual vector to $\mathcal{P}(\Phi_{0})$ has constant angle $\arccos\lambda_{1}$ or $\pi-\arccos\lambda_{1}$ with respect to the unit normal of $\xi\wedge x$ for all $x\in L_{1}$. Namely, $L_{1}^{*}$ is contained in an open hemisphere of $\mathbb{S}^{N_{1}}$ but this cannot happen according to Theorem 5.1.
@@
However, it is unclear to the author at this moment whether one can find a cone over some minimal product (of inputs of positive dimensions and high codimensions) which can support continuous calibrations. If there were no such kind of examples at all, then it indicates that the minimal product structure can automatically produce cones which support no continuous calibrations.
@@ 6. Detect duality obstruction in higher codimension
The second step is to fix a closed smooth submanifold $M$ in $[S]$ which has empty intersection with $S$ and conformally change metric $\hat{g}$ to $h$ in sufficiently small neighborhood of $M$ such that $M$ and $S$ are the only two mass-minimizing in the class $[S]$ with respect to $h$. This can be made rigorously by calibration arguments, for details see [Z12].
@@ 6.2. Generalization to higher codimension
Now we are ready to state what kind of local model $C\subset\mathbb{R}^{N+1}$ can lead to duality obstruction in higher codimension in the last step. Actually there are quite many. The following is a recent configuration result for mass-minimizing cones in Euclidean spaces.
Theorem 6.1 ([Z25]).
@@
Another neat configuration result is Corollary 4.2 of [Z25].
Corollary 6.2.
Let $L$ be an embedded closed minimal subamanifold in $\mathbb{S}^{n}$. Then the cone over minimal product $L\times\mathbb{S}^{d}$ is area-minimizing for sufficiently large $d$.
Both results are proved by Lawlor’s criterion. So our Theorem 1.1 here asserts the existence of coflat calibration forms singular only at the origin for each Lawlor cone in Theorem 6.1 and Corollary 6.2.
@@
This in no way means that the product of any two calibrations is again a calibration. The reason for the conclusion in the quotation is that our singular $(S,o)$ in Example 2 is of codimension one and the coflat calibration has to be a simple form of co-degree one away from the singularity. As a result, in this situation the product of calibrations does produce a coflat calibration with singular set being $X\vee_{(o,o)}X$. Since $(S,o)\times(S,o)$ with singular set $S\vee_{(o,o)}S$ possesses dimension larger than that of $X\vee_{(o,o)}X$, we see that $(S,o)\times(S,o)$ is calibrated by the resulting coflat calibration and the coflat version of fundamental theorem of calibrated geometry now applies.
@@
are calibrated by the coflat calibration $\tilde{\Phi}\wedge\tilde{\Phi}$ in $(X,h)\times(X,h)$. Thus $M\times M$ is homologically mass-minimizing but cannot support any smooth calibration form $\varphi$ globally defined on $(X,h)\times(X,h)$. Otherwise, by the fundamental theorem of calibrated geometry $\varphi$ also calibrates $M\times S$ and similarly it leads to a contradiction with the codimension one of $S$ as follows. Since $\varphi_{(p,o)}$ for $p\in M$ calibrates the tangent cone $C$ at $(p,o)$ - a hyperplane $\times$ hypercone Simons cone $C_{3,3}$, let $\xi$ be the oriented unit top vector of the hyperplane and we get the calibration $\xi\llcorner\varphi_{(p,o)}$ with constant coefficients calibrating $C_{3,3}$. Note that as $C$ splits according to $\mathbb{R}^{8}\oplus\mathbb{R}^{8}$ and the factor $C_{3,3}$ completely sits inside the second $\mathbb{R}^{8}$. So we can observe that the restriction of $\xi\llcorner\varphi_{(p,o)}$ to $\{0\}\times\mathbb{R}^{8}$, again a calibration with constant coefficients but now in $\mathbb{R}^{8}$, calibrates the hypercone $C_{3,3}$.  [^2] If we use $\mathcal{P}$ to stand for the inclusion $\{0\}\times\mathbb{R}^{8}\subset\mathbb{R}^{8}\times\mathbb{R}^{8}$, then the restriction gives exactly $\mathcal{P}^{*}(\xi\llcorner\varphi_{(p,o)})$. This is a contradiction.
@@
in $(X_{1},\tilde{h}_{1})\times(X_{2},\tilde{h}_{2})$. Here $X_{1}\vee_{(p_{i,1},p_{j,2})}X_{2}$ means $X_{1}\times\{p_{j,2}\}\vee_{(p_{i,1},p_{j,2})}\{p_{i,1}\}\times X_{2}$.
@@
A key point is that we modify $\Phi_{i}$ to $\tilde{\Phi}_{i}$ in a small $\epsilon_{i}$-neighborhood $U_{\epsilon_{i}}(M_{i})$ of $M_{i}$ in [Z12] such that $\tilde{\Phi}_{i}$ is a simple form in $U_{3\epsilon_{i}/5}(M_{i})$. So we can multiply a conformal factor function being one in $U_{\epsilon_{i}/5}(M_{i})\cup U_{\epsilon_{i}/5}(S_{i})$ and being a sufficiently large constant in $U_{i}^{c}\triangleq U^{c}_{2\epsilon_{i}/5}(M_{i})\cap U^{c}_{2\epsilon_{i}/5}(S_{i})$ such that under the new metric $\tilde{h}_{i}$ we have the following pointwise
@@
Note that one can choose $\epsilon_{i}$ small enough for $\tilde{\Phi}_{i}$ to be simple in $U_{3\epsilon_{i}/5}(S_{i})\sim\{p_{1,i},p_{2,i}\}$ as well. The reason is that our local model is based on Theorem 1.1 which actually provides coflat calibration singular only at $p_{1,i}$ and $p_{2,i}$ in some neighborhood $\Xi_{i}$ of $S_{i}$. In fact, as on page 10 our local calibration $\varphi$ for a Lawlor cone is simple near $p_{1,i}$ and $p_{2,i}$ and the gluing results in a simple $\Phi_{i}$ in $\Xi_{i}$ away from $p_{1,i}$ and $p_{2,i}$. Therefore, it is easy to see that $\tilde{\Phi}_{1}\wedge\tilde{\Phi}_{2}$ is a calibration in $\big(U_{1}\times X_{2}\big)\cup\big(X_{1}\times U_{2}\big)$ by the simpleness of $\tilde{\Phi}_{i}$ in $U_{3\epsilon_{i}/5}(M_{i})\cup U_{3\epsilon_{i}/5}(S_{i})$.
@@
With respect to the resulting metric $\tilde{h}_{1}\oplus\tilde{h}_{2}$ in Lemma 6.3, the currents in (6.2) are homologically mass-minimzing in the same class and therefore possess the same mass.

### v3

@@ Abstract.
Calibration, comass control, duality obstruction, realization problem
@@ 1. Introduction
In [Z12, z] the author developed a very “soft” framework about extending calibration pairs which play effective roles in detecting duality obstruction and realization problem about singularities (to be explained below). By combining these methods the framework has been carried forward to attack several long-term standing conjectures regarding area-minimizing currents, e.g. [L1, L2, L3].
@@
This obstruction in $(\ast)$ comes from some delicate balance between geometry and topology. And it has been further pointed out on page 431 of [FH] that our obstruction even exists in local around the submanifold in the category of continuous calibration forms. To distinguish this local obstruction $(\ast)$ from global ones, the author used the following example to explain to others since 2020 and such kind of illustration can trace back at least to [FH]. Let $Y=S^{1}\times S^{2020}$ and choose a closed embedded curve $\gamma_{2}$ to be a representative of twice the generator $\tt e$} of $H_{1}(Y;\mathbb{Z})$. Then as in [Z12] we can conformally change background metric such that $\gamma_{2}$ becomes the unique length minimizer in its homology class. As a result, any length minimizer $\gamma_{1}$ in class $\tt e$} cannot be calibrated by any smooth calibration since otherwise the integral current $2\gamma_{1}$ is also calibrated and thus length minimizing in $\gamma_{2}]$ as well which contradicts with the uniqueness in the above. Here the obstruction comes from global in the sense that one cannot apply the idea in a sufficiently small neighborhood of $\gamma_{2}$ for the duality obstruction phenomenon.
In this paper, we extend our (localizable) duality obstruction for question $(\star)$ to genuine high codimension situations beyond that in $(\ast)$. The idea is to find minimizing cones which support coflat calibrations singular only at the origin but cannot be calibrated by any smooth calibration form, and then apply our extension result for calibration pairs in [Z12] (Theorem 3.1 in below).
@@ 2. Basics of calibrations
The dual complex of smooth forms gives us currents. To be more precise, by $(\mathscr{E}^{\prime}_{*}(X),d)$ we mean the dual complex of the de Rham complex of $X$. Elements of $\mathscr{E}^{\prime}_{m}(X)$ are $m$-dimensional de Rham currents (with compact support) and $d$ is the adjoint of exterior differentiation.
@@
A beautiful canonical decomposition lemma for an $m$-form with respect to certain $m$-plane in $\mathbb{R}^{n}$ is the following (its proof in [HL1] inspires a remark for $m=1$ and Lemma 2.12, so we include it in full details).
@@
by $\eta\longmapsto\eta\lrcorner\phi,$ namely, $\lambda(\eta)(v)=\phi(\eta\wedge v)$. Set $\eta_{i}=(-1)^{m+i}v_{1}\wedge\cdots\wedge\hat{v}_{i}\wedge\cdots\wedge v_{m}$ for $i=1,\cdots,m$. Consider forms $\lambda_{i}=\lambda(\eta_{i})$. Then we have
@@ 3. New gluing result for the extension of calibration pairs
where $\tt t$ is a nonzero real number ($|t|\geq 1$ since $\|\phi\|^{*}_{g_{1}}\leq 1$). Hence,
@@
As $\phi$ is not zero, there exists $Q$ in (3.1) which realizes $\|\phi\|^{*2}_{g(s)}=\frac{1}{\|Q\|^{2}}$. Then, by dividing $\tt t^{2}$ on both sides of (3.3) and Lemma 2.2, we get
(3.4) ${\|\phi\|^{*2}_{g(s)}}\leq\frac{1-s}{T(Q;0)}+\frac{s}{T(Q;1)}\leq(1-s)\|\phi\|^{*2}_{g_{1}}+s\|\phi\|^{*2}_{g_{2}}.$
@@
It is worth mentioning that Corollary 3.4 indicates that $\|\phi\|^{*2}_{(\cdot)}$ is a convex functional over the space of metric space. If one uses $F^{a}$ where $a>0$, then we can extend the result accordingly.
Corollary 3.5.
As assumed in the above, we have ${\|\phi\|^{*2a}_{g(s)}}\leq(1-s)\|\phi\|^{*2a}_{g_{1}}+s\|\phi\|^{*2a}_{g_{2}}$.
Proof.
Notice that, for every $a>0$, $F^{a}$ is strictly convex due to the same kind argument. So, for $Q$ in (3.1) realizing $\|\phi\|^{*2}_{g(s)}=\frac{1}{\|Q\|^{2}}$, besides (3.4) we also have
${\|\phi\|^{*2a}_{g(s)}}=\frac{F^{a}(X_{1}(s),\cdots,X_{m}(s))}{{\tt t}^{2a}}\leq\frac{1-s}{T(Q;0)^{a}}+\frac{s}{T(Q;1)^{a}}\leq(1-s)\|\phi\|^{*2a}_{g_{1}}+s\|\phi\|^{*2a}_{g_{2}}.$
Therefore, $\|\phi\|^{*2a}_{(\cdot)}$ is a convex functional over the space of metrics for all $a>0$.
@@ 4. Every Lawlor cone supports calibrations singular only at the origin
In this section, we show that every Lawlor cone can support coflat calibrations which are singular only at the origin of the ambient Euclidean space $(\mathbb{R}^{N+1},g_{E})$. Here by Lawlor cone we mean a minimal cone which can be shown area-minimizing according to Lawlor’s curvature criterion (in practice by the c-control or F-control to be explained below). Such a cone naturally supports non-continuous dual calibrations. Based on some analysis in [z], we are able to construct some coflat calibrations resolving the discontinuity and making the singular set exactly be $\{o\}$. Besides, a shortcut for the realization result in [z] will be explained.
@@
In [Law] Lawlor focused on a similar structure in some suitable angular neighborhood of a minimal cone $C(L^{k})$ instead of in the whole space $\mathbb{R}^{N+1}$. If the boundary of the neighborhood can be mapped to the origin by the projection, then one can map everything outside the neighborhood to the origin. Thus an area-nonincreasing projection is gained and the area-minimality of the cone is proved.
@@
As a result, each Lawlor cone must have some solution $h$ to (4.2) with equality everywhere which touches the $t$-axis at some $t=\tan\theta_{3}>\tan\theta_{0}$ as in the following graph, and moreover, $\theta_{3}<\frac{R(L)}{2}$ as well.
Proof of Theorem 1.1. For point $(t,y)$ with $0<y<\sqrt{t^{2}+1}p(t)$, the inequality (4.2) leads to
@@
Note that $h_{0}$ is strictly deceasing, that is to say $h(t)<p(t)$ in the interval $(0,\tan\theta_{0})$. So, one can arrange $a$ and $\delta$ to make $h_{1}$ strictly decreasing as well. Now one gets $h_{1}(t)<p(t)$ and around the graph of $h_{1}$ in the figure the inequality (4.4) tells that the slope $h^{\prime}(t)$ of a solution to the inequality of (4.2) is allowed to vary from a strictly negative number to a strictly positive number as illustrated in the figure. Then one can deform $h_{1}$ to $h_{2}$ which smoothly joins $1-at^{2}$ and the $t$-axis with its graph touching the axis at $\tan\theta_{2}$ as close as one wishes to the point $\tan\theta_{1}$, and moreover satisfies the calibration inequality (4.2). The reason is that in any small interval $\delta,\epsilon+\delta)$ one can extend $1-at^{2}$ smoothly and make the extension fulfill (4.4) in $(\delta,\epsilon+\delta)$ and (4.5) near $t=\epsilon+\delta$. This allows us to further extend it by (4.5) and it will hit the $t$-axis at a point $\tan\hat{\theta}$ close to $\tan\theta_{1}$. Similarly as what has been done around $t=\delta$, one can deform this solution of inequality (4.2) around $\tan\hat{\theta}$ to make it smoothly tangent to the $t$-axis at $\tan\theta_{2}$ such that the resulting $h_{2}$ also satisfies (4.2) and $\tan\theta_{2}$ can be as close to $\tan\hat{\theta}$ as one wants.
@@
Based on the observation, we know that $C(L_{1}\times\cdots\times L_{n})$ must be calibrated by $\phi_{o}$ which is a calibration with constant coefficients. Fix $x_{i}\in L_{i}$ and let $\xi_{i}$ be the oriented unit tangent $k_{i}$-vector of $T_{x_{i}}L_{i}$ for $i=2,\cdots,n$. Then we shall focus on the calibration $(k_{1}+1)$-form $\Phi_{o}\triangleq\left(\xi_{2}\wedge\cdots\wedge\xi_{n}\right)\lrcorner\phi_{o}$ with constant coefficients. Denote $\mathcal{P}^{*}(\Phi_{o})$ to be the pull-back of $\Phi_{o}$ to the linear subspace $\mathbb{R}^{N_{1}+1}$ via the inclusion $\mathcal{P}:\mathbb{R}^{N_{1}+1}\hookrightarrow\mathbb{R}^{N+1}$ by $\mathcal{P}(x)=(x,0,\cdots,0)$. Note that for $x\in L_{1}$ we have the oriented unit tangent $k_{1}$-vector $\xi=\xi(x)$ of $T_{x}L_{1}$ and $\xi\wedge x$ is the volume form of $C(L_{1})\subset\mathbb{R}^{N_{1}+1}$ (reverse the orientation if need be, here and in the sequel).
Without loss of generality, assume that $L_{1}$ is connected (or restrict ourselves to one connected component of $L_{1}$). We aim to show that evaluation of $\mathcal{P}^{*}(\Phi_{o})$ with $\xi\wedge x$ is a common nonzero constant for all $x\in L_{1}$. First note that evaluation of $\mathcal{P}^{*}(\Phi_{o})$ with $\xi\wedge x$ has the same evaluation value of $\Phi_{o}$ with $\xi\wedge x$. (Here the second $\xi\wedge x$ means the $(k_{1}+1)$-vector $\mathcal{P}_{\#}(\xi\wedge x)$ in $\mathbb{R}^{N+1}$.) Moreover, $\Phi_{o}$ (or $\Phi_{o}$) calibrates all the $(k_{1}+1)$-vectors $\{\xi\wedge p\}$ (short for $\{(\mathcal{P}_{\#}\xi)\wedge p\}$) where $p=p(x)=(\lambda_{1}x,\lambda_{2}x_{2},\cdots,\lambda_{n}x_{n})$ with $x$ running in $L_{1}$. So, for each $x\in L_{1}$, we can extend an orthonormal basis $\{v_{1},\cdots,v_{k_{1}}\}$ of span $\xi$ to an orthonormal basis
@@
(5.4) $\left(\mathcal{P}^{*}(\Phi_{o})\right)(\xi\wedge x)$ $=$ $\Phi_{o}(\xi\wedge x)$
@@
By the assumption, we know that $k_{1}+1=N_{1}$. So $\mathcal{P}^{*}(\Phi_{o})$ is a simple form in $\mathbb{R}^{N_{1}+1}$. In particular, by (5.4) and the connectedness of $L_{1}$ (or its connected component), the fixed unit normal dual vector to $\mathcal{P}^{*}(\Phi_{o})$ has constant angle $\arccos\lambda_{1}$ or $\pi-\arccos\lambda_{1}$ with respect to the unit normal of $\xi\wedge x$ for all $x\in L_{1}$. Namely, $L_{1}^{*}$ is contained in an open hemisphere of $\mathbb{S}^{N_{1}}$ but this cannot happen according to Theorem 5.1.
@@
However, it is unclear to the author at this moment whether one can find a cone over some minimal product (of inputs of positive dimensions) which can support continuous calibrations. If there were no such kind of examples at all, then it indicates that the minimal product structure can automatically produce cones which support no continuous calibrations.
@@ 6. Detect duality obstruction in higher codimension
The second step is to fix a closed smooth submanifold $M$ in $[S]$ which has empty intersection with $S$ and conformally change metric $\hat{g}$ to $h$ in sufficiently small neighborhood of $M$ such that $M$ and $S$ are the only two mass-minimizing representatives in the class $[S]$ with respect to $h$. This can be made rigorously by calibration arguments, for details see [Z12].
@@ 6.2. Generalization to higher codimension
Now we are ready to state what kind of local model $C\subset\mathbb{R}^{N+1}$ can lead to duality obstruction in higher codimension in the last step. Actually there are quite many. The following is a recent general configuration result for mass-minimizing cones in Euclidean spaces.
Theorem 6.1 (Theorem 1.1 of [Z25]).
@@
Another neat configuration result is the following.
Theorem 6.2 (Theorem 1.3 of [Z25]).
Let $L$ be an embedded closed minimal subamanifold in $\mathbb{S}^{n}$. Then the cone over minimal product $L\times\mathbb{S}^{d}$ in $\mathbb{R}^{n+d+2}$ is area-minimizing when $d$ is sufficiently large.
Both results are proved by Lawlor’s criterion. So our Theorem 1.1 here asserts the existence of coflat calibration forms singular only at the origin for each Lawlor cone in Theorem 6.1 and Theorem 6.2.
@@
This in no way means that the Cartesian product of any two calibrations is again a calibration. In fact a recent result of [Zust25] shows the opposite and thus gives a negative answer to a long-standing Federer question (of comass version). The reason for the conclusion in the quotation is that our singular $(S,o)$ in Example 2 is of codimension one and the coflat calibration has to be a simple form of co-degree one away from the singularity. As a result, in this situation the product of calibrations does produce a coflat calibration with singular set being $X\vee_{(o,o)}X$. Since $(S,o)\times(S,o)$ with singular set $S\vee_{(o,o)}S$ possesses dimension larger than that of $X\vee_{(o,o)}X$, we see that $(S,o)\times(S,o)$ is calibrated by the resulting coflat calibration and the coflat version of fundamental theorem of calibrated geometry now applies.
@@
are calibrated by the coflat calibration $\tilde{\Phi}\wedge\tilde{\Phi}$ in $(X,h)\times(X,h)$. Thus $M\times M$ is homologically mass-minimizing but cannot support any smooth calibration form $\varphi$ globally defined on $(X,h)\times(X,h)$. Otherwise, by the fundamental theorem of calibrated geometry $\varphi$ also calibrates $M\times S$ and similarly it leads to a contradiction with the codimension one of $S$ as follows. Since $\varphi_{(p,o)}$ for $p\in M$ calibrates the tangent cone $C$ at $(p,o)$ - a hyperplane $\times$ hypercone Simons cone $C_{3,3}$, let $\xi$ be the oriented unit top vector of the hyperplane and we get the calibration $\xi\lrcorner\varphi_{(p,o)}$ with constant coefficients calibrating $C_{3,3}$. Note that as $C$ splits according to $\mathbb{R}^{8}\oplus\mathbb{R}^{8}$ and the factor $C_{3,3}$ completely sits inside the second $\mathbb{R}^{8}$. So we can observe that the restriction of $\xi\lrcorner\varphi_{(p,o)}$ to $\{0\}\times\mathbb{R}^{8}$, again a calibration with constant coefficients but now in $\mathbb{R}^{8}$, calibrates the hypercone $C_{3,3}$.  [^2] If we use $\mathcal{P}$ to stand for the inclusion $\{0\}\times\mathbb{R}^{8}\subset\mathbb{R}^{8}\times\mathbb{R}^{8}$, then the restriction gives exactly $\mathcal{P}^{*}(\xi\lrcorner\varphi_{(p,o)})$. This is a contradiction.
@@
in $(X_{1},\tilde{h}_{1})\times(X_{2},\tilde{h}_{2})$. Here $X_{1}\vee_{(p_{i,1},p_{j,2})}X_{2}$ means $\big(X_{1}\times\{p_{j,2}\}\big)\cup\big(\{p_{i,1}\}\times X_{2}\big)$.
@@
Recall that we modify $\Phi_{i}$ to $\tilde{\Phi}_{i}$ in a small $\epsilon_{i}$-neighborhood $U_{\epsilon_{i}}(M_{i})$ of $M_{i}$ in [Z12] such that $\tilde{\Phi}_{i}$ is a simple form in $U_{3\epsilon_{i}/5}(M_{i})$. So we can multiply a conformal factor function being one in $U_{\epsilon_{i}/5}(M_{i})\cup U_{\epsilon_{i}/5}(S_{i})$ and being a sufficiently large constant in $U_{i}^{c}\triangleq U^{c}_{2\epsilon_{i}/5}(M_{i})\cap U^{c}_{2\epsilon_{i}/5}(S_{i})$ such that under the new metric $\tilde{h}_{i}$ we have the following pointwise
@@
Note that one can choose $\epsilon_{i}$ small enough for $\tilde{\Phi}_{i}$ to be simple in $U_{3\epsilon_{i}/5}(S_{i})\sim\{p_{1,i},p_{2,i}\}$ as well. The reason is that our local model is based on Theorem 1.1 which actually provides coflat calibration singular only at $p_{1,i}$ and $p_{2,i}$ in some neighborhood $\Xi_{i}$ of $S_{i}$. In fact, as in §3.2 our local calibration $\varphi$ for a Lawlor cone is simple near $p_{1,i}$ and $p_{2,i}$ and the gluing results in a simple $\Phi_{i}$ in $\Xi_{i}$ away from $p_{1,i}$ and $p_{2,i}$. Therefore, it is easy to see that $\tilde{\Phi}_{1}\wedge\tilde{\Phi}_{2}$ is a calibration in ${\mathcal{S}}^{c}\cap\Big(\big(U_{1}\times X_{2}\big)\cup\big(X_{1}\times U_{2}\big)\Big)$ by the simpleness of $\tilde{\Phi}_{i}$ in $U_{3\epsilon_{i}/5}(M_{i})\cup U_{3\epsilon_{i}/5}(S_{i})$.
@@
With respect to the resulting metric $\tilde{h}_{1}\oplus\tilde{h}_{2}$ in Lemma 6.3, the currents in (6.2) are homologically mass-minimzing in the same class.